# On dense subalgebras of the singular ideal in groupoid C*-algebras

### v1

@@ 2.1. Characterisation of density
Let $I=\overline{J\cap\mathscr{C}_{c}(G)}$. Since $\mathscr{C}_{c}(G)$ is dense in $C^{*}_{r}(G)$, $I$ is an ideal and $I\subseteq J$. By [16, Theorem 5.5 & Proposition 5.19], we have $\eta_{x}(J\cap\mathscr{C}_{c}(G))=J_{x}\cap\mathbb{C}[G^{x}_{x}]$. Continuity of $\eta_{x}$ then implies $J_{x}\cap\mathbb{C}[G^{x}_{x}]\subseteq\eta_{x}(I)\subseteq\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$. Since $\eta_{x}(I)=I_{x}$ is closed, it follows that $I_{x}=\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$. Therefore, Theorem 2.5 shows that $I=J$ if and only if $\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}=J_{x}$ for all $x\in G^{(0)}$, as desired. ∎
@@
The “only if” direction of Proposition 2.6 holds for all étale groupoids $G$ because $\eta_{x}(J\cap\mathscr{C}_{c}(G))\subseteq J_{x}\cap\mathbb{C}[G^{x}_{x}]$ holds trivially ($J_{x}$ here means the isotropy fibre from the reduced groupoid $C^{*}$-algebra). It can be shown using the calculation of isotropy fibres of the singular ideal in [16, Theorem 5.5] that the first example in [20] of a non-Hausdorff groupoid $G$ with $J\cap\mathscr{C}_{c}(G)$ not dense in $J$ has an isotropy fibre $J_{\epsilon}\neq\{0\}$ with $J_{\epsilon}\cap\mathbb{C}[\Gamma]=\{0\}$. Therefore, this remark may be applied to this example to provide an alternate proof of non-density.
Following [16], we now explain how the density question can be reduced to one about explicit ideals in group $C^{*}$-algebras. Let $\Gamma$ be an amenable discrete group. Let $\text{Sub}(\Gamma)$ denote the set of all subgroups in $\Gamma$. Consider the map
@@
where $\mathbbm{1}_{X}\colon\Gamma\to\{0,1\}$ is the characteristic function, and equip $\{0,1\}^{\Gamma}$ with the product topology. We equip $\text{Sub}(\Gamma)$ with the topology induced by this map, known as the Chabauty topology, making $\text{Sub}(\Gamma)$ a Stone space (this topology agrees with the Fell topology of [14]).
For a subgroup $X\in\text{Sub}(\Gamma)$, write $\Gamma/X$ for the set of left-cosets. Let $\lambda_{\Gamma/X}\colon\Gamma\to B\big(\ell^{2}(\Gamma/X)\big)$ denote the associated quasi-regular representation as described in [3]. Concretely, we have
$\lambda_{\Gamma/X}(g)\delta_{hX}\coloneq\delta_{ghX}$ for $g\in\Gamma$ and $hX\in\Gamma/X$, where $\delta_{hX}\in\ell^{2}(\Gamma/X)$ denotes a standard basis vector. We also write $\lambda_{\Gamma/X}$ for the associated -representation on $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$.
Let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be closed in the Chabauty topology and invariant under conjugation by elements of $\Gamma$. Following [16], we let the ideal $J_{\Gamma,\mathcal{X}}$ denote the $C^{*}$-kernel of $\lambda_{\Gamma/\mathcal{X}}\coloneq\oplus_{X\in\mathcal{X}}\lambda_{\Gamma/X}$ inside $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$. Concretely, we have $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\text{ker}(\lambda_{\Gamma/X})$.
@@
Let $G$ be an amenable second-countable étale groupoid. For $x\in G^{(0)}$ write $\mathcal{X}(x)\coloneq\pi^{-1}(x)\cap\tilde{G}^{(0)}_{{\rm{ess}}}$, where $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is the canonical surjection and $\tilde{G}^{(0)}_{\rm{ess}}$ is the unit space of the essential Hausdorff cover (see Subsection 1.2). The isotropy groups of $G$ are amenable, and in this case it was shown in [16, Definition 5.1 & Theorem 5.5] that $J_{x}=J_{G^{x}_{x},\mathcal{X}(x)}$ for every $x\in G^{(0)}$. The following Theorem is then immediate.
@@
Since $J_{\Gamma,\mathcal{X}}=\{0\}$ whenever $\{e\}\in\mathcal{X}$, in the above characterisation one only needs to check whether $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ whenever $\{x\}\notin\mathcal{X}(x)$. These units $x$ are shown in [16, Corollary 5.6] to be precisely the extremely dangerous points described in [24, Proposition 1.12]. Note also that the set of units $x\in G^{(0)}$ for which $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ is a $G$-invariant set.
@@
Assume that there exists a (discrete) countable amenable group $\Gamma$ which fails to have the Density Property, and let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be a closed conjugation-invariant set of subgroups for which $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is not dense in $J_{\Gamma,\mathcal{X}}$, i.e. $\mathcal{X}\notin\mathcal{D}_{\Gamma}$. Note that the trivial subgroup, $\{e\}$, is not in $\mathcal{X}$ (amenability of $\Gamma$ ensures that the quasi-regular representation $\lambda_{\Gamma/\{e\}}$ has trivial kernel). By [16, Section 6], there exists an amenable second-countable étale groupoid $G$ and a unit element $x_{0}\in G^{(0)}$ for which $G^{x_{0}}_{x_{0}}=\Gamma$ and $\mathcal{X}(x_{0})=\mathcal{X}$. It follows from Theorem 2.9 that $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. Therefore, we have established the following.
@@
The intersection $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for all amenable second-countable étale groupoids if and only if all countable amenable groups satisfy the Density Property.
@@ 3. Groups with the Density Property
In this section, we study the set $\mathcal{D}_{\Gamma}$ for amenable discrete groups $\Gamma$, and show that certain classes of groups satisfy the Density Property (see Definition 2.8). Using Theorem 2.9, we are then able to deduce that $J\cap\mathscr{C}_{c}(G)$ is dense in the singular ideal $J$ for certain classes of groupoids.
Throughout this section we work with the reduced group $C^{*}$-algebra $C^{*}_{r}(\Gamma)$ since this is canonically isomorphic to the full $C^{*}$-algebra. The set of subgroups $\text{Sub}(\Gamma)$ will always be equipped with the Chabauty topology, and given a closed and conjugation-invariant subset $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ we write $J_{\Gamma,\mathcal{X}}\coloneq\bigcap_{X\in\mathcal{X}}\text{ker}\lambda_{\Gamma/X}\trianglelefteq C^{*}_{r}(\Gamma)$ (see Subsection 2.1).
Given a subgroup $\Lambda\subseteq\Gamma$, we write $E_{\Lambda}\colon C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Lambda)$ for the canonical conditional expectation and $I_{\Lambda}\colon C^{*}_{r}(\Lambda)\to C^{*}_{r}(\Gamma)$ for the canonical inclusion. If $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ is closed and conjugation-invariant, then $\Lambda\cap\mathcal{X}\coloneq\{\Lambda\cap X:X\in\mathcal{X}\}$ is a closed and conjugation-invariant set of subgroups of $\Lambda$ by [16, Proposition 7.4]. We isolate the following observations from [16, Propositions 7.4 & 7.7].
@@
    Let $N$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. Then,
    $E_{N}(J_{\Gamma,\mathcal{X}})=J_{N,\mathcal{X}}\quad\text{and}\quad E_{N}(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma])=J_{N,\mathcal{X}}\cap\mathbb{C}[N].$

### v2

@@ 3. Characterisation of density
Let $I\coloneq\overline{J\cap\mathscr{C}_{c}(G)}$. Since $\mathscr{C}_{c}(G)$ is dense in $C^{*}_{r}(G)$, $I$ is an ideal and $I\subseteq J$. By [19, Theorem 5.5 & Proposition 5.19] we have $\eta_{x}(J\cap\mathscr{C}_{c}(G))=J_{x}\cap\mathbb{C}[G^{x}_{x}]$. Continuity of $\eta_{x}$ then implies $J_{x}\cap\mathbb{C}[G^{x}_{x}]\subseteq\eta_{x}(I)\subseteq\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$. Since $\eta_{x}(I)=I_{x}$ is closed, it follows that $I_{x}=\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$. Therefore, Theorem 2.5 shows that $I=J$ if and only if $\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}=J_{x}$ for all $x\in G^{(0)}$. ∎
@@
The “only if” direction of Proposition 3.1 holds for all étale groupoids $G$ because $\eta_{x}(J\cap\mathscr{C}_{c}(G))\subseteq J_{x}\cap\mathbb{C}[G^{x}_{x}]$ holds trivially ($J_{x}$ here means the isotropy fibre from the reduced groupoid $C^{*}$-algebra). It can be shown using the calculation of isotropy fibres of the singular ideal in [19, Theorem 5.5] that the first example in [25] of a non-Hausdorff groupoid $G$ with $J\cap\mathscr{C}_{c}(G)$ not dense in $J$ has an isotropy fibre $J_{\epsilon}\neq\{0\}$ with $J_{\epsilon}\cap\mathbb{C}[\Gamma]=\{0\}$. Therefore, this remark may be applied to this example to provide an alternate proof of non-density.
Following [19], we now explain how the density question can be reduced to one about explicit ideals in group $C^{*}$-algebras. Let $\Gamma$ be an amenable discrete group. Let $\text{Sub}(\Gamma)$ denote the set of all subgroups of $\Gamma$. Consider the map
@@
where $\mathbbm{1}_{X}\colon\Gamma\to\{0,1\}$ is the indicator function, and equip $\{0,1\}^{\Gamma}$ with the product topology. We equip $\text{Sub}(\Gamma)$ with the topology induced by this map, known as the Chabauty topology, making $\text{Sub}(\Gamma)$ a Stone space (this topology agrees with the Fell topology of [14]).
For a subgroup $X\in\text{Sub}(\Gamma)$, write $\Gamma/X$ for the set of left-cosets. Let $\lambda_{\Gamma/X}\colon\Gamma\to B\big(\ell^{2}(\Gamma/X)\big)$ denote the associated quasi-regular representation as described in [4]. Concretely, we have $\lambda_{\Gamma/X}(g)\delta_{hX}\coloneq\delta_{ghX}$ for $g\in\Gamma$ and $hX\in\Gamma/X$, where $\delta_{hX}\in\ell^{2}(\Gamma/X)$ denotes a standard basis vector. We also write $\lambda_{\Gamma/X}$ for the associated -representation on $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$.
Let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$. Following [19], we let the ideal $J_{\Gamma,\mathcal{X}}$ denote the intersection $\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})$ inside $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$. We will primarily be interested in the case where $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ is closed in the Chabauty topology and invariant under conjugation by elements of $\Gamma$.
@@
Let $G$ be an amenable second-countable étale groupoid. For $x\in G^{(0)}$, recall from Definition 1.3 that $\mathcal{X}(x)\coloneq\pi^{-1}\big(\{x\}\big)\cap\tilde{G}^{(0)}_{{\rm{ess}}}$ is a closed conjugation-invariant collection of subgroups of the isotropy group $G^{x}_{x}$. The isotropy groups of $G$ are amenable, and in this case [19, Definition 5.1 & Theorem 5.5] showed that $J_{x}=J_{G^{x}_{x},\mathcal{X}(x)}$ for every $x\in G^{(0)}$. The following theorem is then immediate.
@@
In Theorem 3.4 it suffices to check whether $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ for extremely dangerous points $x\in G^{(0)}$ (as described in [30, Proposition 1.12]). This is because $J_{\Gamma,\mathcal{X}}=\{0\}$ whenever the trivial group $\{e\}$ belongs to $\mathcal{X}$, and a unit $x\in G^{(0)}$ satisfies $\{x\}\notin\mathcal{X}(x)$ if and only if it is extremely dangerous (see [19, Corollary 5.6]).
@@
Assume that there exists a (discrete) countable amenable group $\Gamma$ that fails to have the Density Property, and let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be a closed conjugation-invariant set of subgroups for which $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is not dense in $J_{\Gamma,\mathcal{X}}$ (i.e. $\mathcal{X}\notin\mathcal{D}_{\Gamma}$). In particular, the trivial subgroup $\{e\}$ does not belong to $\mathcal{X}$. By [19, Section 6], there exists an amenable and second-countable étale groupoid $G$ and a unit element $x_{0}\in G^{(0)}$ for which $G^{x_{0}}_{x_{0}}=\Gamma$ and $\mathcal{X}(x_{0})=\mathcal{X}$. It follows from Theorem 3.4 that $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. Therefore, we have established the following.
@@
The intersection $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for all second-countable amenable étale groupoids if and only if all (discrete) countable amenable groups satisfy the Density Property.
@@ 4. Groups with the Density Property
In this section, we study the set $\mathcal{D}_{\Gamma}$ for amenable discrete groups $\Gamma$, and show that certain classes of groups satisfy the Density Property (see Definition 3.3). We do this in three parts. In Subsection 4.1 we simplify the conditions one needs to check for finding elements of the set $\mathcal{D}_{\Gamma}$. In Subsection 4.2 we prove some permanence results for the Density Property. We then use these observations in Subsection 4.3 to prove Theorem 4.16, which in particular states that all countable finite-by-nilpotent groups satisfy the Density Property. Combining with Theorem 3.4, we are able to deduce that $J\cap\mathscr{C}_{c}(G)$ is dense in the singular ideal $J$ for certain classes of groupoids.
Throughout the section we work with the reduced group $C^{*}$-algebra $C^{*}_{r}(\Gamma)$ since this is canonically isomorphic to the full $C^{*}$-algebra. The set of subgroups $\text{Sub}(\Gamma)$ will always be equipped with the Chabauty topology, and given a closed and conjugation-invariant subset $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ we write $J_{\Gamma,\mathcal{X}}\coloneq\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})\trianglelefteq C^{*}_{r}(\Gamma)$ (see Section 3).
Let us record three preliminary lemmas which will be useful throughout the subsections. For a subgroup $\Lambda\subseteq\Gamma$, we write $I_{\Lambda}\colon C^{*}_{r}(\Lambda)\to C^{*}_{r}(\Gamma)$ for the canonical inclusion and $E_{\Lambda}\colon C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Lambda)$ for the canonical conditional expectation. By [19, Proposition 7.4], $\Lambda\cap\mathcal{X}\coloneq\{\Lambda\cap X\colon X\in\mathcal{X}\}$ is a closed and conjugation-invariant set of subgroups of $\Lambda$. In the first lemma we isolate the following observations from [19, Propositions 7.4 & 7.7] which describe the images of $J_{\Lambda,\Lambda\cap\mathcal{X}}$ and $J_{\Gamma,\mathcal{X}}$ under the inclusion and conditional expectation maps.
@@
    Let $N\subseteq\Gamma$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. Then,
    $E_{N}\big(J_{\Gamma,\mathcal{X}}\big)=J_{N,\mathcal{X}}\quad\text{and}\quad E_{N}\big(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]\big)=J_{N,\mathcal{X}}\cap\mathbb{C}[N].$